OpenStudy (anonymous):

how do you simplify this? please help (3x^3y^4)^2

OpenStudy (anonymous):

exponents get distributed too \[(3x^3y^4)^2=3^2x^{3^2}y^{4^2}\]

OpenStudy (campbell_st):

ummm not quite right the power of a power laws says multiply the powers \[(x^a)^b = x^{a \times b}\] so to simplify its \[3^2 x^{ 3 \times 2} y ^{ 4 \times 2}\] simplify the powers ans evaluate 3^2